The potential difference between a certain thundercloud and the ground is 7MVolts. Find the energy dissipated when a charge of 50 Coul is transferred from the clouds to the ground in a lightning stroke



1
Expert's answer
2022-06-14T03:45:01-0400

Given:

"V=7*10^6\:\rm V"

"q=50\:\rm C"


The dissipated energy is given by

"W=qV=50*7*10^6=3.5*10^8\:\rm J"